OpenStudy (anonymous):

Find the mean of this set of numbers. 9, –6, –8, 11, –4, 10

OpenStudy (cwrw238):

add them and then divide by 6

OpenStudy (anonymous):

okay thanks

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

are you happy adding negative numbers eg 2 + -6 ?

OpenStudy (anonymous):

-8

OpenStudy (cwrw238):

no -2 + -6 = -8 2 + -6 = 2 - 6 = -4

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

To find mean you add all the numbers and divide by how many numbers there are.There are six numbers, so you divide by 6. So all these numbers added equals 12. 12 divided by 6 equals 2. The answer is 2. Understand?

OpenStudy (anonymous):

ohh can u help me on a different question??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Identify which property of equality is used to transform Equation 1 to Equation 2. Equation 1: 5 + x = 9 Equation 2: x = 4

OpenStudy (cwrw238):

yes - thiese can be tricky maybe the analogy of a thermometer will help